The pKb of the cyanide ion (CN-) is 4.60. What is the pOH and the pH of a 0.05 M cyanide solution?
Answer #1

CN- + H2O ---> HCN + OH-

Kb = [HCN][OH-]/[CN-]

pKb = 4.60

-logKb = 4.60

Kb = antilog(-4.60) = 2.51*10-5

Kb = [HCN][OH-]/[CN-]

2.51*10-5 = x.x/0.05

x2 = 1.25*10-6M

x = (1.25*10-6)1/2 = 1.12*10-3M = [OH-]

pOH = -logOH- = -log(1.12*10-3) = 2.95

pH*pOH = Kw

pH = Kw/pOH

= 10-14/2.95 = 3.38*10-15

pH = 3.38*10-15